OpenStudy (anonymous):

What is the vertex of the graph of y = 5(x + 4)2 + 3? (−4, 3) (4, 3) (3, 4) (3, −4)

OpenStudy (onepieceftw):

Look at the vertex formula y= a(x-h)^2+k vertex = (h,k) In your equation h=-4 and k=3

OpenStudy (elise_a18):

Simplify the expression then use the formula for slope: y=mx+b. B is the y-intercept and m is the slope.\[\frac{ rise }{ run}\]

OpenStudy (anonymous):

So the first one?